# A Large Recurrent Action Model: xLSTM enables Fast Inference for Robotics Tasks

### v1

@@ 4.2 Scaling comparison
To conduct our main comparisons, we train our four backbone variants on the full training task mixture of 432 tasks. For each architecture backbone, we report performance scores for four model sizes: 16M, 48M, 108M, and 206M parameters. We train all models for 200K updates with a batch size of 128 and context length of 50 timesteps. All domains are represented with approximately equal proportion, resulting in 33K updates per domain. Additional implementation details and hyperparameters for every backbone variant and model size are available in Appendix B.
Sequence prediction performance. In Figure 2a, we report the validation set perplexity for all backbones and model sizes averaged over the individual scores from all domains. To achieve this, we maintain a hold-out set of trajectories for each training task (2.5%) and compute the perplexities after every 50K steps. Both recurrent backbones outperform the Transformer baseline considerably, especially as the model sizes increase. We provide the perplexities on the training set in Figure 13.
Evaluation performance. During training, we evaluate our agents after every 50K step in all 432 training environments. In Figure 2b, we report the resulting normalized performances averaged across all six domains. The recurrent backbones outperform the Transformer one across model sizes. While xLSTM and Mamba performs similarly at smaller scales, xLSTM tends to outperform Mamba at larger scales (206M). This is an important advantage of xLSTM, as LRAM agents can strongly benefit from more data and consequently larger models. Note, that Mamba has a significantly higher number of parameters than competitors. For the zero-shot evaluation performances on the 37 hold-out tasks, we refer to Figure 15 in Appendix C.2.
Performance per domain. In Figure 3, we report the normalized scores for the 206M parameter models attained on all six domains. For Meta-World, DMControl, Mimicgen, Composuite, and Procgen we use data-normalized scores, as suggested by [Levine et al.,, 2020]. For Atari, we report human-normalized scores. Overall, we observe that the xLSTM backbone outperforms competitors on three of the six domains, while all methods perform similarly on the remaining 3 domains.
These experiments suggest that modern recurrent backbones can be attractive alternatives to the Transformer architecture for building LAMs in terms of final performance.
@@ 4.3 Analyses & Ablations
Fine-tuning. To assess the effect of the recurrent backbones on fine-tuning performance, we fine-tune our models on 37 held-out environments from all 6 domains. We evaluate the fine-tuning performance of the xLSTM architecture for both the 16M parameter pretrained models and compared it against an xLSTM trained from scratch. The pretrained LRAM outperforms the randomly initialized xLSTM model in most domains. For detailed results, see Appendix C.3. This suggests that fine-tuning performance is not negatively affected by switching the backbone.
Figure 4: ICL with modern recurrent architectures on Dark-Room $10\times 10$.
In-context Learning. Next, we study the ICL abilities of our recurrent backbones on the Dark-Room environment considered in prior work on in-context RL [Laskin et al.,, 2022; Lee et al.,, 2023; Schmied et al., 2024b, ]. To study ICL in isolation, we train models from scratch with a multi-episodic context, which results in a large context length (we refer to Appendix C.4 for details on the experiment setup). In particular, we adopt the Algorithm Distillation (AD, Laskin et al.,, 2022) framework and exchange the Transformer backbone architecture with modern recurrent architectures. In Figure 17, we report the ICL performance on (a) 80 train and (b) 20 hold-out tasks. We find that xLSTM [7:1] attains the highest overall scores both on training and hold-out tasks, which we attribute to the state-tracking abilities [Merrill et al.,, 2024] of sLSTM blocks.
Embedding space analysis. In Figure 5, we analyze the representations learned by our model. To this end, we sample 32 sub-trajectories from every task, extract the sequence representation at the last layer, cluster them using UMAP [McInnes et al.,, 2018], and color every point by its domain. Appendix E describes the setup in greater detail. We find that tasks from the same domain cluster together. Furthermore, xLSTM exhibits a more refined domain separation compared to DT, which may contribute to the better down-stream performance.
@@
Figure 5: Embedding space comparison. UMAP clustering of hidden states for all tasks for 16M models, colored by domain. xLSTM exhibits a better domain separation than DT.
Removing Actions & Effect of Context Length. We found that removing actions from the context results in better performance across backbones. While context lengths beyond 1 hurt performance on Meta-World when training with actions, the reverse is true when training without actions (see Figure 23). This is in contrast to recent works, which did not benefit from longer contexts [Octo Model Team et al.,, 2024]. While removing actions improves performance on the robotics domains, it does not affect performance on discrete control. For robotics, we observed that the models become overly confident (high action logits), which is problematic if poor initial actions are produced. We assume this is because in robotics actions change smoothly and by observing previous actions the agent learns shortcuts. Removing actions from the input prevents the agent from using shortcuts. Importantly, the evaluation performance improves as the sequence length increases, which indicates that the history helps to predict the next action (e.g., by observing mistakes made in the recent past).
We present additional ablations on the effect of reducing the number of layers in xLSTM and disabling Dropout on DT in Appendix D.3 and D.2, respectively.
@@ 4.4 Inference Time Comparison
Finally, we empirically examine the difference between xLSTM-based and Transformer-based agents at inference time. Similar to De et al., [2024], we report both latency and throughput. We focus our analysis on latency, as it is the more important dimension for real-time applications.
(a) $B=1$
(b) $B=16$
Figure 6: Latency comparison on A100. We report latency for varying context lengths (in timesteps) with fixed batch sizes $B$ of 1 and 16. We compare DT to xLSTM with the same number of layer blocks and parameters on Atari Freeway. Missing bars for DT indicate out-of-memory (OOM).
Setup. We conduct all inference time tests on A100 GPUs with 40GB of RAM using 206M parameter models. For the Transformer, we use KV-caching and FlashAttention [Dao,, 2023] as supported by PyTorch [Paszke et al.,, 2019]. For xLSTM, we use recurrent-style inference using custom kernels to accelerate the computations (see Figure 22 for the impact of kernel acceleration). For both backbones, we use torch.compile. The Transformer with KV-caching has a linear time complexity per step and quadratic in the sequence length. In contrast, the xLSTM has a constant time complexity per step and linear in the sequence length. Therefore, we expect speed-ups especially for longer sequences and larger batch sizes, as observed by De et al., [2024]. To ensure a fair comparison, we compare DT and xLSTM with the same number of layer blocks and increase the hidden size of xLSTM to match the number of parameters of DT (see Appendix D.3 for evaluation performance of these models). We provide further details on our inference time tests in Appendix C.5.
@@
Figure 7: Memory consumption during Latency comparison on A100 (of GPU memory) for varying context lengths and $B=1$.
Latency. Similar to De et al., [2024], we measure latency by the average time (in seconds) taken to perform a single inference step with a fixed batch size $B$ (lower is better). In Figure, 6, we report the latencies for varying context lengths, $C\in[50,25600]$ and two batch sizes $B\in\{1,16\}$. Note that $C$ is in time steps and every time step contains 3 tokens (state, reward-to-go, reward). Hence, the effective sequence length for the largest $C$ is 76800. As expected, we find that the recurrent backbone attains lower inference latencies than the Transformer one. As the sequence length increases, DT runs out of memory due to the increasing size of the KV cache (see Figure 7). In contrast, the inference speeds for xLSTM are independent of the context length, and therefore enable significantly longer context lengths. This property is particularly interesting for in-context RL, which requires keeping multiple episodes in the context [Laskin et al.,, 2022]. Nevertheless, our experiments highlight that the materialization of the complexity advantage (quadratic vs. linear) depends on the device, model size, batch size and the context length, which is similar to findings by De et al., [2024].
Throughput. Throughput is measured by the total amount of inference steps performed per second for a model with a fixed context length. In Figure, 8, we report the throughputs for varying batch sizes, $B\in[1,128]$ for a fixed context length of $C=1600$. Here, the batch size can be interpreted as the number of parallel environments the agent interacts with. As expected, we find that xLSTM attains considerably higher throughputs than the DT. The benefit of xLSTM increases with larger batch sizes. While the DT with quadratic complexity in the sequence length goes OOM for batch sizes above 64, the xLSTM with linear complexity can easily handle larger batch sizes. In both experiments, the recurrent xLSTM performs favorably over the Transformer backbone.
@@ 5 Conclusion
In this work, we study the aptitude of modern recurrent architectures as alternatives to Transformers for building LAMs. We found that our LRAM with an xLSTM or Mamba at its core compare favorably to the Transformer in terms of evaluation performance across different model scales. Moreover, we demonstrated that xLSTM-based LRAMs exhibit higher inference speeds, especially at large context sizes. Thus, the empirical evidence suggests, that recurrent backbones such as the xLSTM can be attractive alternatives for LAMs. Notably, the linear-time inference complexity of xLSTM may enable applications that require long context lengths, such as in-context RL, and facilitate the application of large-scale agents for real-time applications, such as robotics.
Limitations. The primary target application of LAMs is robotics. While the majority of our experiments involve robotic simulations, we do not yet provide empirical evidence for real robots. We do, however, believe that our findings translate to real-world scenarios and aim to provide further evidence in future work. Moreover, the fine-tuning experiments in this work are limited to offline RL. We envision that an agent pre-trained by behavioral cloning on large-scale offline RL datasets may be successfully fine-tuned in an online RL setting to explore new strategies that do not appear in the training data. Modern recurrent architectures offer both parallel and recurrent training mode, which might be the key to success for such applications. While we provide initial evidence of improved ICL abilities of modern recurrent architectures, we only consider a limited grid-world setting. Consequently, we aim to further investigate the in-context RL abilities of recurrent backbones on more complex environments in future work.
6 Ethics Statement
@@
7 Reproducibility
Our code-base used for our experiments and the datasets we generated are publicly available at: https://github.com/ml-jku/LRAM. We describe the environments we use for our experiments and provide dataset statistics in Appendix A. Furthermore, in Appendix B, we provide implementation details for all methods and a list of hyperparameters used for our experiments. In Appendix C, we present additional figures that accompany our results in the main text (e.g., all model sizes). Finally, in Appendices D and E, we provide further details on the conducted ablation studies and the embedding space analysis, respectively.
Acknowledgments and Disclosure of Funding
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, and Leonardo at CINECA, Italy. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.

### v2

@@ 3.2 Large Recurrent Action Models (LRAMs)
4 Experiments
We study the aptitude of modern recurrent architectures as LAMs on 432 tasks from 6 domains: Atari (Bellemare et al., 2013), Composuite (Mendez et al., 2022), DMControl (Tassa et al., 2018), Meta-World (Yu et al., 2020b), Mimicgen (Mandlekar et al., 2023), and Procgen (Cobbe et al., 2020b). To this end, we compile a large-scale dataset containing 894 million transitions (see Section 4.1). Across all experiments, we compare four backbone variants: xLSTM [7:1], xLSTM [1:0] (Beck et al., 2024), Mamba (Gu & Dao, 2023), and the GPT-2 style Transformer employed in the DT (Chen et al., 2021). Following (Beck et al., 2024), we use the bracket notation for xLSTM, which indicates the ratio of mLSTM to sLSTM blocks. For example, xLSTM [1:0] contains only mLSTM blocks.
In Section 4.2, we conduct a scaling comparison for four model sizes ranging from 16M to 208M parameters that shows that modern recurrent architectures achieve performance comparable or favorable to the Transformer baseline across different model sizes. In Section 4.3, we study the impact of the recurrent backbones on fine-tuning performance, ICL abilities, and further analyze our trained recurrent backbones. Finally, in Section 4.4, we empirically examine the differences at inference time in terms of latency and throughput between xLSTM and Transformer-based agents, which indicate advantages for the recurrent backbone.
4.1 Datasets & Environments
Datasets. We compile a large-scale dataset comprising 432 tasks from six domains. We leverage datasets from prior works if they exist. For Atari, we extract 5M transitions per task from the DQN-Replay dataset released by Agarwal et al. (2020). For Composuite, we leverage the datasets released by (Hussing et al., 2023). For Meta-World, we use 2M transitions per task released by (Schmied et al., 2024a). For DMControl, we generate 10M transitions per task using task-specific RL agents. For Mimicgen, we use the datasets for the 21 tasks released by (Mandlekar et al., 2023) and generate trajectories for the remaining 62 tasks. Finally, for Procgen, we extract 20M transitions from the datasets released by (Schmied et al., 2024b). Our final dataset contains 3.4M trajectories and in total 894M transitions (see Table 1). We reserve an additional 37 tasks from the same domains for zero-shot evaluation. To foster future research, we release our data-preparation pipeline and generated data.
Environments. Atari and Procgen come with image observations and discrete action. In contrast, the remaining four domains exhibit state-based observations and continuous actions. Consequently, our experiments involve a mixture of state and action spaces as well as varying episode lengths (see Table 1). Periodically evaluating the trained agents on all 432 tasks sequentially is time-consuming and we, therefore, distributed the evaluation across GPUs and parallel processes (see Appendix C). Additional details on our datasets and environments are available in Appendix B.
@@ 4.2 Scaling comparison
To conduct our main comparisons, we train our four backbone variants on the full training task mixture of 432 tasks. For each architecture backbone, we report performance scores for four model sizes: 16M, 48M, 108M, and 206M parameters. We train all models for 200K updates with a batch size of 128 and context length of 50 timesteps. All domains are represented with approximately equal proportion, resulting in 33K updates per domain. Additional implementation details and hyperparameters for every backbone variant and model size are available in Appendix C.
Sequence prediction performance. In Figure 2a, we report the validation set perplexity for all backbones and model sizes averaged over the individual scores from all domains. To achieve this, we maintain a hold-out set of trajectories for each training task (2.5%) and compute the perplexities after every 50K steps (see Figure 12 for training perplexities). Both recurrent backbones outperform the Transformer baseline considerably, especially as the model sizes increase.
Evaluation performance. During training, we evaluate our agents after every 50K step in all 432 training environments. In Figure 2b, we report the resulting normalized performances averaged across all six domains. The recurrent backbones outperform the Transformer one across model sizes. While xLSTM and Mamba performs similarly at smaller scales, xLSTM tends to outperform Mamba at larger scales (206M). This is an important advantage of xLSTM, as LRAM agents can strongly benefit from more data and consequently larger models. Note, that Mamba has a significantly higher number of parameters than competitors. For the zero-shot evaluation performances on the 37 hold-out tasks, we refer to Figure 14 in Appendix D.2.
Performance per domain. In Figure 3, we report the normalized scores for the 206M models attained on all six domains. For Meta-World, DMControl, Mimicgen, Composuite, and Procgen we use data-normalized scores, as suggested by (Levine et al., 2020). For Atari, we report human-normalized scores. We observe that the xLSTM backbone outperforms competitors on three of the six domains, while they perform similarly on the remaining domains.
@@ 4.3 Analyses & Ablations
Fine-tuning. To assess the effect of the recurrent backbones on fine-tuning performance, we fine-tune our models on 37 held-out environments from all 6 domains. We evaluate the fine-tuning performance of the xLSTM architecture for the 16M pretrained models and compare it against an xLSTM trained from scratch. The pretrained LRAM outperforms the randomly initialized xLSTM model in most domains (see Figure 15). This suggests that fine-tuning performance is not affected negatively by switching the backbone.
Figure 4: In-context Learning with modern recurrent architectures on 20 hold-out tasks for Dark-Room $10\times 10$.
In-context Learning. Next, we study the ICL abilities of our recurrent backbones on the Dark-Room environment considered in prior work on in-context RL (Laskin et al., 2022; Lee et al., 2023; Schmied et al., 2024b). To study ICL in isolation, we train models from scratch with a multi-episodic context, which results in a large context length (see Appendix D.4 for details on the experiment setup). In particular, we adopt the Algorithm Distillation (AD, Laskin et al., 2022) framework and exchange the Transformer backbone architecture with modern recurrent architectures. In Figure 4, we report the ICL performance on the 20 hold-out tasks (see Figure 16 for training tasks). We find that xLSTM [7:1] attains the highest overall scores both on the 80 training and 20 hold-out tasks, which we attribute to the state-tracking abilities (Merrill et al., 2024) of sLSTM blocks.
Embedding space analysis. In Figure 5, we analyze the representations learned by our model. We sample 32 sub-trajectories from every task, extract the sequence representation at the last layer, cluster them using UMAP (McInnes et al., 2018), and color every point by its domain (see Appendix F for more details). We find that tasks from the same domain cluster together. Furthermore, xLSTM exhibits a more refined domain separation compared to DT, which may further contribute to the better down-stream performance.
@@
Figure 5: Embedding space comparison. UMAP clustering of hidden states for all tasks for 16M, colored by domain. xLSTM exhibits a better domain separation than DT.
(a) Latency, $B=1$
(b) Latency, $B=16$
(c) Memory, $B=1$
Figure 6: Latency comparison on A100. We report latency for varying context lengths (in timesteps) with batch sizes (a) $B=1$ and (b) $B=16$. In (c), we show the memory consumption in % of GPU memory with $B=1$. We compare DT to xLSTM with the same number of layer blocks and parameters on Atari Freeway. Missing bars for DT indicate out-of-memory (OOM).
Removing Actions & Effect of Context Length. We found that removing actions from the context results in better performance across backbones. While context lengths beyond 1 hurt performance on Meta-World and DMControl and when training with actions, the reverse is true when training without actions (see Figures 22, 23, 25). This is in contrast to recent works, which did not benefit from longer contexts (Octo Model Team et al., 2024). While removing actions improves performance on Meta-World/DMControl, it does not affect performance on discrete control environments. For Meta-World/DMControl, we observed that the models become overly confident, which is problematic if poor initial actions are produced. This is because many robotics environments exhibit smoothly changing actions and by observing previous actions the agent learns shortcuts. A similar issue has been observed by Wen et al. (2020) and termed the copycat problem. Removing actions from the input prevents the agent from using shortcuts and alleviates the copycat problem. Importantly, the evaluation performance improves across domains as the sequence length increases, which indicates that the history helps to predict the next action (e.g., by observing mistakes made in the past, see Figures 24, 26).
Return-conditioning vs. Behavior Cloning. Across our experiments, we utilized a sequence representation that includes return-to-go tokens, as commonly used in DTs (Chen et al., 2021; Lee et al., 2022). However, many recent works focus on behavior cloning without return conditioning (Reed et al., 2022; Brohan et al., 2023a). Therefore, we study the effect of excluding the RTG/reward tokens from the sequence at the 206M parameter scale, to validate that our findings transfer to the behavior cloning setting. Indeed, we find that the same trends hold (see Figures 27 and 28).
mLSTM-to-sLSTM Ratio. Throughout experiments, we compare two xLSTM variants: xLSTM [7:1] and xLSTM [1:0]. These ratios were proposed by Beck et al. (2024) and we maintain the same ratios for consistency (see Appendix C.3). While mLSTM is parallelizable, sLSTM enables state-tracking (Merrill et al., 2024). To better understand the effect of the ratio, we conduct ablation studies both on the 432 tasks and on Dark-Room (see Appendix E.3), similar to Beck et al. (2024). We find that other ratios, such as [3:1], can be effective, and highlight the importance of placing sLSTMs at lower-level layers (Figure 30). However, the effectiveness of sLSTM layers is dependent on the task at hand. Complex tasks with long horizons or partial observability, as are common in real-world applications, may benefit from the state-tracking abilities provided by sLSTM.
We present additional ablations on the effect of reducing the number of layers in xLSTM and disabling Dropout on DT in Appendix E.5 and E.4, respectively.
@@ 4.4 Inference Time Comparison
Finally, we empirically examine the difference between xLSTM-based and Transformer-based agents at inference time. Similar to De et al. (2024), we report both latency and throughput. We focus our analysis on latency, as it is the more important dimension for real-time applications.
Setup. We conduct all inference time tests on A100s with 40GB of RAM using 206M models. For the Transformer, we use KV-caching and FlashAttention (Dao, 2023) as supported by PyTorch (Paszke et al., 2019). For xLSTM, we use recurrent-style inference using custom kernels to accelerate computations (see Figure 21 for the impact of kernel acceleration). For both backbones, we use torch.compile. The Transformer with KV-caching has a linear time complexity per step and quadratic in the sequence length. In contrast, the xLSTM has a constant time complexity per step and linear in the sequence length. Therefore, we expect speed-ups especially for longer sequences and larger batch sizes, as observed by De et al. (2024). To ensure a fair comparison, we compare DT and xLSTM with the same number of layer blocks and increase the hidden size of xLSTM to match the number of parameters of DT (see Appendix E.5 for evaluation performance of these models). We provide further details on our inference time tests in Appendix D.5.
@@
Latency. Similar to De et al. (2024), we measure latency by the average time (in seconds) taken to perform a single inference step with a fixed batch size $B$ (lower is better). In Figure, 6, we report the latencies for varying context lengths, $C\in[50,25600]$ and two batch sizes $B\in\{1,16\}$. Note that $C$ is in time steps and every time step contains 3 tokens (state, reward-to-go, reward). Hence, the effective sequence length for the largest $C$ is 76800. As expected, we find that the recurrent backbone attains lower inference latencies than the Transformer one. As the sequence length increases, DT runs out of memory due to the increasing size of the KV cache (see Figure 6c). In contrast, the inference speeds for xLSTM are independent of the context length, and therefore enable significantly longer context lengths. This property is particularly interesting for in-context RL, which requires keeping multiple episodes in the context (Laskin et al., 2022). Nevertheless, our experiments highlight that the materialization of the complexity advantage depends on the device, model size, batch size and the context length, which is similar to findings by De et al. (2024).
@@
Throughput. Throughput is measured by the total amount of inference steps performed per second for a model with a fixed context length. In Figure, 7, we report the throughputs for varying batch sizes, $B\in[1,128]$ for a fixed context length of $C=1600$. Here, the batch size can be interpreted as the number of parallel environments the agent interacts with. As expected, we find that xLSTM attains considerably higher throughputs than the DT. The benefit of xLSTM increases with larger batch sizes. While the DT with quadratic complexity in the sequence length goes OOM for batch sizes above 64, the xLSTM with linear complexity can easily handle larger batch sizes. In both experiments, the recurrent xLSTM performs favorably over the Transformer backbone.
@@ 5 Conclusion
In this work, we study the aptitude of modern recurrent architectures as alternatives to Transformers for building LAMs. We found that our LRAM with an xLSTM or Mamba at its core compare favorably to the Transformer in terms of evaluation performance across model scales (see Section 4.2). Moreover, we demonstrated that xLSTM-based LRAMs exhibit higher inference speeds, especially at large context sizes (see Section 4.4). Thus, the empirical evidence suggests, that recurrent backbones can be attractive alternatives for LAMs. Notably, the linear-time inference complexity of xLSTM may enable applications that require long context lengths (e.g., ICL), and facilitate the application of large-scale agents for real-time applications, such as robotics.
Modern recurrent architectures and Transformers come with different pros and cons. xLSTM and Mamba, on the one hand, exhibit a complexity advantage over Transformers. Their linear complexity ensures that the computational requirements increase slower with the sequence length, which enables more efficient inference and is particularly relevant for edge-applications. While we conduct our inference time comparisons on a high-end data-center GPU, applications on edge-devices may have to deal with less powerful accelerators. Importantly, we found that LAMs strongly benefit from longer sequences (see Section 4.3). Transformers, on the other hand, are effective for applications that require exact recall of tokens in a sequence, which can be important for decision-making (Ni et al., 2024). Finally, xLSTM in particular enables state-tracking via sLSTM blocks, which Transformers and Mamba cannot perform (Merrill et al., 2024). State tracking can be important for logic tasks or dealing with partial observability and may be a useful tool for practitioners. Given these differences, different backbones should be considered depending on the task at hand.
Limitations. The primary target application of LAMs is robotics. While the majority of our experiments involve robotic simulations, we do not yet provide experiments for real robots. We do, however, believe that our findings translate to real-world scenarios and aim to provide further evidence in future work. Moreover, our fine-tuning experiments are limited to offline RL. We envision that an agent pre-trained on large-scale datasets can be successfully fine-tuned via online RL to explore new strategies that do not appear in the training data. Modern recurrent architectures offer both parallel and recurrent training mode, which might be the key to success for such applications. While we provide evidence for improved ICL abilities of LRAM, we only consider a grid-world setting. We aim to further investigate the ICL abilities of LRAM in more complex environments.
Impact Statement
@@
Acknowledgements
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, and Leonardo at CINECA, Italy. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/COE12]. We thank the projects INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.